OpenStudy (anonymous):

A car starting from rest moves with constant acceleration of 2.0 m/s2 for 10 s, then travels with constant speed for another 10 s, and then finally slows to a stop with constant acceleration of -2.0 m/s2. How far does it travel? the options are 200 m 400 m 300 m 500 m can you explain what equation to use and show the steps???? thanks so much!

OpenStudy (shamim):

s= ut+(1/2) at^2

OpenStudy (shamim):

v= u+ at

OpenStudy (shamim):

s= vt

OpenStudy (shamim):

u can use these equations

OpenStudy (gtxmuqsit):

40 meters I guess

OpenStudy (anonymous):

from what we know, which of the 3 kiematic equations can we use to solve? we will want to pick an equation that a, v, t in it becuase those are what the problem gives us And we'll want to pick the equation that has s in it also. s stands for position or distance, which is what we are looking for, so the s should be the only thing on = side of the equation. v=v+at v^2=v^2+2as s = s + vt + (1/2)at^2 the method of how we will find how far the car has travled will be to find out how far the car travels form time0 to time1. then, we'll repeat that again and find how far the car travels from time1 to time2. then repeat again from time2 to time3.